# Zitterbewegung, momentum and spin dynamics of electromagnetic wave in linear dielectric medium

### v3

@@ Abstract
The momentum of light in dielectric media has been a century-long controversy that continues to attract significant interest. In a linear dielectric medium with refractive index n, the momentum is predicted to be smaller by a factor of n according to Max Abraham, and larger by the same factor according to Hermann Minkowski. By studying the coupled dynamics of electromagnetic waves and dipoles in a dielectric medium, we show that the change in momentum of the dipole, expressed by the Lorentz force, corresponds to the Abraham momentum and is given by the expectation value of the spin-projected momentum vector. On the other hand, the Minkowski momentum is obtained as the magnitude of the spin-projected momentum vector from the energy-momentum dispersion relation derived by diagonalizing the coupled Hamiltonian and determines the direction of refraction in accordance with Snell’s law. Our model also predicts a zitterbewegung-like oscillation due to helicity mixing between left- and right-handed wave components, mediated by dipole oscillation. These internal wave dynamics may be observable via wavepacket motion or polarization-sensitive measurements.
@@ I Introduction
Due to the opposing predictions by Max Abraham and Hermann Minkowski, the momentum of light in dielectric media has been a century-long debate [1, 2, 3]. Since the wavelength in a dielectric medium with refractive index $n$ is shortened, Minkowski predicted that the wavevector $k=2\pi n/\lambda$ and momentum are larger:
$p_{\mathrm{M}}=\hbar k=\frac{nhc}{\lambda}.$ (1)
@@
$p_{\mathrm{A}}=\frac{\hbar\omega}{c^{2}}v=\frac{hc}{n\lambda}.$ (2)
@@
The rest of this article is organized as follows. Section II introduces the Drude dipole model and derives the refractive index. Section III formulates the Hamiltonian matrix describing electromagnetic wave dynamics and presents the helicity mixing between left- and right-handed components mediated by dipole coupling. Section IV presents the main results, including the derivation of Minkowski and Abraham momenta, velocity operator, spin precession, and a zitterbewegung-like oscillation. Section V discusses connections to previous models and possible experimental consequences, and Section VI concludes. The emergence of zitterbewegung-like behavior from internal mode coupling in the medium highlights wave dynamics that resemble features of particle models, bridging wave and particle pictures.
@@ II.1 Drude dipole model
The classical Drude model focuses on the dynamics of electrons under an electric field $\mathbf{E}$. Here, we include the dynamics of the medium by modeling the positive charge as connected to the electron by a spring with natural oscillation frequency $\omega_{0}$, yielding coupled equations:
@@
where $m_{-}$ and $m_{+}\gg m_{-}$ are the masses of negative and positive charges, respectively, and $m=(m_{+}^{-1}+m_{-}^{-1})^{-1}$ is the effective mass of the relative dipole motion. The first term on the right-hand side represents the internal restoring force of the dipole, with natural oscillation frequency $\omega_{0}$. $\mathbf{B}$ is the magnetic field.
@@
Equation 5 is the Drude equation for an electron attached to a spring with natural oscillation frequency $\omega_{0}$. For $m\ll M$, $\dot{\mathbf{R}}$ can be neglected. Using Fourier transform $f(x,t)=(2\pi)^{-1}\int e^{i(kx-\omega t)}f(\omega)d\omega$, the equation can be linearized and solved:
@@
Note that while the initial derivation uses a single dipole oscillator for conceptual clarity, all dynamical quantities in the model-such as energy, polarization, and velocity-are defined per unit volume. This is implemented by dividing by the representative volume $V$, so that the resulting dynamics describe the averaged response of a polarization mode within the dielectric.
@@
where $\omega_{p}=\sqrt{e^{2}/(\varepsilon_{0}mV)}$ is the plasma frequency, $1/V$ is the density, and $\varepsilon_{0}$ is the vacuum permittivity. In the low-frequency regime, the system behaves as a linear medium with refractive index
@@
The derivation of this refractive index using a classical oscillator model is well known and appears in standard textbooks [21]. Ref. [26] discusses similar dipole dynamics in the context of laser cooling. Compared to that treatment, we neglect spatial derivatives of the electric and magnetic fields $\nabla\textbf{E}$ and $\nabla\textbf{B}$, and focus on homogeneous field interactions in momentum space.
@@ II.2 Equation of motion of light
Coupled equations of dipole and EM wave arise from the Drude equation and Maxwell’s equations for current $\mathbf{j}=\dot{\mathbf{P}}=-e\dot{\mathbf{r}}_{e}/V$:
@@
Here, $\hat{\mathbf{p}}=-i\hbar\nabla$, and $\mathbf{s}$ is a Hermitian SO(3) rotation representation [27] fulfilling the angular momentum commutation relations $s_{i},s_{j}]=i\sum_{k}\epsilon_{ijk}s_{k}$ for spin-1, with $s^{2}=2$, and $s_{z}=0,\pm 1$.
@@
representing the right-handed and left-handed wave functions of EM wave [28], and
@@
$H_{c}$ $=\left[\begin{array}[]{cc}c\hat{\mathbf{p}}\cdot\mathbf{s}&0\\ 0&-c\hat{\mathbf{p}}\cdot\mathbf{s}\end{array}\right],\quad H_{a}=\left[\begin{array}[]{cc}\hbar\omega_{0}&0\\ 0&-\hbar\omega_{0}\end{array}\right],$ (17)
@@
$H_{ac}$ enables indirect coupling between right- and left-handed EM waves. Since $H$ is Hermitian, the Heisenberg equation can be used to find time derivatives: $df/dt=[f,H]/(i\hbar)$.
@@ III Result
The spin dynamics satisfy
@@
$\langle\hat{\mathbf{v}}\rangle=\Psi^{\dagger}\hat{\mathbf{v}}\Psi=c\left(\mathbf{c}_{+}^{\dagger}\mathbf{s}\mathbf{c}_{+}-\mathbf{c}_{-}^{\dagger}\mathbf{s}\mathbf{c}_{-}\right).$ (27)
@@ III.1 Minkowski momentum from diagonalization
Here, $n$ is the refractive index from Eq. 8. The dispersion relations, corresponding to eigenfrequencies of the coupled system, are shown in Fig. 2. The first element $\hat{H}_{11}$ of the diagonalized matrix corresponds to the EM wave dispersion in the linear medium, yielding the Minkowski momentum
$p_{M}=\left<\hat{\textbf{p}}\cdot\textbf{s}\right>=\frac{n}{c}\left<\hat{H}_{11}\right$>, (33)
@@
Figure 2: Dispersion in a linear dielectric medium: hybridization of right-handed (blue dashed line), left-handed EM waves (red dashed line), and dipole (green dotted line) modes results in linear dispersion $\hbar\omega=\pm pc/n$ (purple lines) at low frequencies.
@@
Figure 3: (a) Refraction conserves tangential momentum $\hat{\mathbf{p}}_{t}$ at the interface. The velocity of EM wave $\hat{\mathbf{v}}$ in the medium is a superposition of right-handed ($\mathbf{c}_{+}$) and left-handed ($\mathbf{c}_{-}$) EM waves, whose velocities are parallel and anti-parallel to the spin $\mathbf{s}$, respectively. The spin $\mathbf{s}$ precesses around $\hat{\mathbf{p}}$ with $s_{z}\parallel\hat{\mathbf{p}}$. (b) In metamaterials, the left-handed EM wave ($\mathbf{c}_{-}$) contribution dominates, resulting in velocity opposite to $\hat{\mathbf{p}}$.
@@ III.2 Abraham momentum from acceleration
Thus far, we have identified conserved quantities of the EM wave in the medium: the momentum operator $\hat{p}$, the spin component $s_{z}$ parallel to it, and energy. The EM wave energy density entering the medium equals to the sum of electromagnetic and dipole oscillation energy densities,
@@
$\frac{d\langle\hat{\mathbf{v}}\rangle}{dt}=\frac{e\dot{\mathbf{r}}_{e}\times c\mathbf{B}}{\frac{V\varepsilon_{0}}{2}\left(E^{2}+c^{2}B^{2}\right)+\frac{1}{2}\left(m\omega_{0}^{2}r_{e}^{2}+m\dot{r}_{e}^{2}\right)}.$ (45)
Here, the expectation value of the first term is zero because $\left<\textbf{s}\right>\parallel\hat{\textbf{p}}$. This relates acceleration to the reaction force on the dipole Lorentz force in Eq. 4:
@@
Here, $\{\hat{\mathbf{p}}\cdot\mathbf{s},\mathbf{s}\}/2$ denotes the component of $\hat{\mathbf{p}}$ parallel to the spin. Taking the expectation value of Eq. 51 yields additional terms beyond Eq. 46:
@@ IV Discussion
The novelty of our model lies in unifying spin dynamics, momentum transfer, and polarization precession within a semiclassical first-quantization framework. The two forms of electromagnetic momentum arise from different aspects of spin-projected operators. The Abraham momentum is given by the expectation value of the spin-projected momentum vector $\{\hat{\mathbf{p}}\cdot\mathbf{s},\mathbf{s}\}/2$ (Eq. 57), directly linking it to the Lorentz-force reaction on the medium. In contrast, the Minkowski momentum corresponds to the magnitude $p_{M}=\langle\hat{\mathbf{p}}\cdot\mathbf{s}\rangle$ (Eq. 33), obtained from the dispersion relation of the diagonalized Hamiltonian. This coexistence is reminiscent of the velocity of a Dirac relativistic electron: its magnitude equals the speed of light, yet its expectation value is reduced due to zitterbewegung. This operator-level distinction provides a unified microscopic interpretation of how both momenta coexist and fulfil complementary physical roles in dielectric media.
@@
In Ref. [31], the photon mass drag and light-matter momentum transfer are computed using continuum field theory and relativistic energy-momentum conservation. While the resulting Abraham momentum is consistent with ours, their approach does not consider the spin dynamics of the EM wave. Our Hamiltonian allows one to track the coupled evolution of spin, velocity and linear momentum simultaneously, yielding a natural operator-level distinction between the two momenta.
In Ref. [32], the spin and orbital angular momentum (OAM) of light are investigated using canonical field decomposition. While their work thoroughly characterizes the transfer and conservation of optical spin and OAM, they embeds the effect of medium on the frequency-dependent permittivity $\varepsilon(\omega)$. In contrast, our model explicitly includes dipole degrees of freedom without absorbing these effects into an effective bulk parameter. This permits a more detailed examination of how polarization mediates the coupling between right- and left-handed EM modes.
The coupled dynamics of the EM wave modes in our model closely resemble the zitterbewegung predicted for relativistic electrons in the Dirac equation. Recent experimental work supports the physical plausibility of zitterbewegung-like dynamics in photonic systems. Ref. [33] reported transverse oscillations in the position of light wavepackets confined in planar and lattice microcavities. These oscillations were interpreted as a manifestation of zitterbewegung, arising from interference between two polarization-dependent optical modes. Although the experiment did not measure helicity or spin directly, the internal interference mechanism provides a useful analogy to the zitterbewegung predicted in our model. This lends conceptual support to the idea that wave-packet dynamics governed by coupled internal degrees of freedom can lead to observable oscillatory effects in optical systems.
@@ V Conclusion
Acknowledgments.
We acknowledge the support of Universitas Indonesia via PUTI Research Grant PKS-182/UN2.RST/HKP.05.00/2025

### v4

# Zitterbewegung, momentum and spin dynamics of electromagnetic waves in linear dielectric medium

@@ Abstract
The momentum of light in dielectric media has been a century-long controversy that continues to attract significant interest. In a linear dielectric medium with refractive index n, the momentum is predicted to be smaller by a factor of n according to Abraham, and larger by the same factor according to Minkowski. By studying the coupled dynamics of electromagnetic waves and dipoles in a dielectric medium, we show that the change in momentum of the dipole, expressed by the Lorentz force, corresponds to the Abraham momentum and is given by the expectation value of the spin-projected momentum vector. On the other hand, the Minkowski momentum is obtained as the magnitude of the spin-projected momentum vector from the energy-momentum dispersion relation derived by diagonalizing the coupled Hamiltonian and determines the direction of refraction in accordance with Snell’s law. Our model also predicts a zitterbewegung-like oscillation due to helicity mixing between left- and right-handed wave components, mediated by dipole oscillation. These internal wave dynamics may be observable via wavepacket motion or polarization-sensitive measurements.
@@ I Introduction
Due to the opposing predictions by Abraham and Minkowski, the momentum of light in dielectric media has been a century-long debate [1, 2, 3]. Since the wavelength in a dielectric medium with refractive index $n$ is shortened, Minkowski predicted that the wavevector $k=2\pi n/\lambda$ and momentum are larger:
$p_{M}=\hbar k=\frac{nhc}{\lambda}.$ (1)
@@
$p_{A}=\frac{\hbar\omega}{c^{2}}v=\frac{hc}{n\lambda}.$ (2)
@@
The rest of this article is organized as follows. Section II.1 introduces the Drude dipole model and derives the refractive index. Section II.2 formulates the Hamiltonian matrix describing electromagnetic wave dynamics and presents the helicity mixing between left- and right-handed components mediated by dipole coupling. Section III presents the main results, including the derivation of Minkowski and Abraham momenta, velocity operator, spin precession, and a zitterbewegung-like oscillation. Section IV discusses connections to previous models and possible experimental consequences. Section V summarizes. The emergence of zitterbewegung-like behavior from internal mode coupling in the medium highlights wave dynamics that resemble features of particle models, bridging wave and particle pictures.
@@ II.1 Drude dipole model
The classical Drude model focuses on the dynamics of electrons under an electric field $\mathbf{E}$. Here we include the dynamics of the medium by modeling the positive charge as connected to the electron by a spring with natural oscillation frequency $\omega_{0}$, yielding coupled equations
@@
where $m_{-}$ and $m_{+}\gg m_{-}$ are the masses of negative and positive charges, respectively, and $m=(m_{+}^{-1}+m_{-}^{-1})^{-1}$ is the effective mass of the relative dipole motion. The first term on the right-hand side represents the internal restoring force of the dipole, with natural oscillation frequency $\omega_{0}$, and $\mathbf{B}$ is the magnetic field.
@@
Equation 5 is the Drude equation for an electron attached to a spring with natural oscillation frequency $\omega_{0}$. For $m\ll M$, $\dot{\mathbf{R}}$ can be neglected. Using the Fourier transform $f(x,t)=(2\pi)^{-1}\int e^{i(kx-\omega t)}f(\omega)d\omega$, the equation can be linearized and solved:
@@
Note that while the initial derivation uses a single dipole oscillator for conceptual clarity, all dynamical quantities in the model, such as energy, polarization, and velocity, are defined per unit volume. This is implemented by dividing by the representative volume $V$ so that the resulting dynamics describe the averaged response of a polarization mode within the dielectric.
@@
where $\omega_{p}=\sqrt{e^{2}/\varepsilon_{0}mV}$ is the plasma frequency, $1/V$ is the density, and $\varepsilon_{0}$ is the vacuum permittivity. In the low-frequency regime, the system behaves as a linear medium with refractive index
@@
The derivation of this refractive index using a classical oscillator model is well known and appears in standard textbooks [21]. Reference [26] discusses similar dipole dynamics in the context of laser cooling. Compared to that treatment, we neglect spatial derivatives of the electric and magnetic fields $\nabla\textbf{E}$ and $\nabla\textbf{B}$, and focus on homogeneous field interactions in momentum space.
@@ II.2 Equation of motion of light
Coupled equations of the dipole and the EM wave arise from the Drude equation and Maxwell’s equations for current $\mathbf{j}=\dot{\mathbf{P}}=-e\dot{\mathbf{r}}_{e}/V$:
@@
Here $\hat{\mathbf{p}}=-i\hbar\nabla$ and $\mathbf{s}$ is a Hermitian SO(3) rotation representation [27] fulfilling the angular momentum commutation relations $s_{i},s_{j}]=i\sum_{k}\epsilon_{ijk}s_{k}$ for spin 1, with $s^{2}=2$, and $s_{z}=0,\pm 1$.
@@
representing the right-handed and left-handed wave functions of the EM wave [28] and
@@
$H_{c}$ $=\left[\begin{array}[]{cc}c\hat{\mathbf{p}}\cdot\mathbf{s}&0\\ 0&-c\hat{\mathbf{p}}\cdot\mathbf{s}\end{array}\right],\quad H_{a}=\left[\begin{array}[]{cc}\hbar\omega_{0}&0\\ 0&-\hbar\omega_{0}\end{array}\right]$ (17)
@@
The $H_{ac}$ enables indirect coupling between right- and left-handed EM waves. Since $H$ is Hermitian, the Heisenberg equation can be used to find time derivatives: $df/dt=[f,H]/i\hbar$.
@@ III Result
The spin dynamics satisfies
@@
$\langle\hat{\mathbf{v}}\rangle=\Psi^{\dagger}\hat{\mathbf{v}}\Psi=c(\mathbf{c}_{+}^{\dagger}\mathbf{s}\mathbf{c}_{+}-\mathbf{c}_{-}^{\dagger}\mathbf{s}\mathbf{c}_{-}).$ (27)
@@ III.1 Minkowski momentum from diagonalization
Here $n$ is the refractive index from Eq. 8. The dispersion relations, corresponding to eigenfrequencies of the coupled system, are shown in Fig. 2. The first element $\hat{H}_{11}$ of the diagonalized matrix corresponds to the EM wave dispersion in the linear medium, yielding the Minkowski momentum
$p_{M}=\left<\hat{\textbf{p}}\cdot\textbf{s}\right>=\frac{n}{c}\langle\hat{H}_{11}\rangle,$ (33)
@@
Figure 2: Dispersion in a linear dielectric medium. Hybridization of right-handed (blue dashed line), left-handed EM waves (red dashed line), and dipole (green dashed line) modes results in linear dispersion $\hbar\omega=\pm pc/n$ (purple solid lines) at low frequencies.
@@
Figure 3: (a) Refraction conserves tangential momentum $\hat{\mathbf{p}}_{t}$ at the interface. The velocity of the EM wave $\hat{\mathbf{v}}$ in the medium is a superposition of right-handed ($\mathbf{c}_{+}$) and left-handed ($\mathbf{c}_{-}$) EM waves, whose velocities are parallel and anti-parallel to the spin $\mathbf{s}$, respectively. The spin $\mathbf{s}$ precesses around $\hat{\mathbf{p}}$ with $s_{z}\parallel\hat{\mathbf{p}}$. (b) In metamaterials, the left-handed EM wave $\mathbf{c}_{-}$ contribution dominates, resulting in velocity opposite to $\hat{\mathbf{p}}$.
@@ III.2 Abraham momentum from acceleration
Thus far, we have identified conserved quantities of the EM wave in the medium: the momentum operator $\hat{p}$, the spin component $s_{z}$ parallel to it, and energy. The EM wave energy density entering the medium equals the sum of electromagnetic and dipole oscillation energy densities,
@@
$\frac{d\langle\hat{\mathbf{v}}\rangle}{dt}=\frac{e\dot{\mathbf{r}}_{e}\times c^{2}\mathbf{B}}{\frac{V\varepsilon_{0}}{2}\left(E^{2}+c^{2}B^{2}\right)+\frac{1}{2}\left(m\omega_{0}^{2}r_{e}^{2}+m\dot{r}_{e}^{2}\right)}.$ (45)
Here the expectation value of the first term is zero because $\left<\textbf{s}\right>\parallel\hat{\textbf{p}}$. This relates acceleration to the reaction force on the dipole Lorentz force in Eq. 4:
@@
Taking the expectation value of Eq. 51 yields additional terms beyond Eq. 46:
@@
Here $\hat{\mathbf{p}}_{\mathbf{s}}$ denotes the component of $\hat{\mathbf{p}}$ parallel to the spin.
$\hat{\textbf{p}}_{\textbf{s}}=\left[\begin{array}[]{cccc}\frac{1}{2}\{\hat{\textbf{p}}\cdot\mathbf{s},\mathbf{s}\}&0&0&0\\ 0&\frac{1}{2}\{\hat{\textbf{p}}\cdot\mathbf{s},\mathbf{s}\}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right].$
@@ IV Discussion
The novelty of our model lies in unifying spin dynamics, momentum transfer, and polarization precession within a semiclassical first-quantization framework. The two forms of electromagnetic momentum arise from different aspects of spin-projected operators. The Abraham momentum is related to the expectation value of the spin-projected momentum vector $\{\hat{\textbf{p}}\cdot\mathbf{s},\mathbf{s}\}/2$, directly linking it to the Lorentz-force reaction on the medium (Eq. 52). In contrast, the Minkowski momentum corresponds to the magnitude $p_{M}=\langle\hat{\mathbf{p}}\cdot\mathbf{s}\rangle$ (Eq. 33), obtained from the dispersion relation of the diagonalized Hamiltonian. This coexistence is reminiscent of the velocity of a Dirac relativistic electron: its magnitude equals the speed of light, yet its expectation value is reduced due to zitterbewegung. This operator-level distinction provides a unified microscopic interpretation of how both momenta coexist and fulfill complementary physical roles in dielectric media.
@@
In Ref. [31] the photon mass drag and light-matter momentum transfer are computed using continuum field theory and relativistic energy-momentum conservation. While the resulting Abraham momentum is consistent with ours, their approach does not consider the spin dynamics of the EM wave. Our Hamiltonian allows one to track the coupled evolution of spin, velocity and linear momentum simultaneously, yielding a natural operator-level distinction between the two momenta.
In Ref. [32] the spin and orbital angular momentum (OAM) of light are investigated using canonical field decomposition. While their work thoroughly characterizes the transfer and conservation of optical spin and OAM, they embed the effect of the medium on the frequency-dependent permittivity $\varepsilon(\omega)$. In contrast, our model explicitly includes dipole degrees of freedom without absorbing these effects into an effective bulk parameter. This permits a more detailed examination of how polarization mediates the coupling between right- and left-handed EM modes.
The coupled dynamics of the EM wave modes in our model closely resemble the zitterbewegung predicted for relativistic electrons in the Dirac equation. Recent experimental work supports the physical plausibility of zitterbewegung-like dynamics in photonic systems. Reference [33] reported transverse oscillations in the position of light wavepackets confined in planar and lattice microcavities. These oscillations were interpreted as a manifestation of zitterbewegung, arising from interference between two polarization-dependent optical modes. Although the experiment did not measure helicity or spin directly, the internal interference mechanism provides a useful analogy to the zitterbewegung predicted in our model. This lends conceptual support to the idea that the wave-packet dynamics governed by coupled internal degrees of freedom can lead to observable oscillatory effects in optical systems.
@@ V Conclusion
Acknowledgment
We acknowledge supports from Universitas Indonesia via PUTI Research Grant No. PKS-182/UN2.RST/HKP.05.00/2025